Please help me solve this equation:x-1/x+1-6/x-3=3

Answer by geetha_rama(94)

(X-1)/(X+1) - 6/X - 3 =
(X-1)/(X+1) - 6/X = 6
tAKING lcm
{X(X-1) - 6(X+1)}/X(X+1) = 6
(X^2 - X -6X -6 )/(X^2 + X) = 6
X^2 - X -6X -6 = 6* (X^2 + X)
X^2 - X -6X -6 = 6X^2 + 6X
=> X^2 - 7X - 6 = 6X^2 + 6X
5X^2 -13X +6 =0
5X^2 -10X - 3X -6 = 0
5X(X-2) -3(X-2)=0
(X-2)(5X-3) = 0
=> X = 2 OR 3/5